OpenStudy (anonymous):

Given f(x) = 2x^2 + 2x and g(x) = x + 1, find the function (f/g)(x).

OpenStudy (shamil98):

(f/g)(x) = divide the values of f(x) by the values of g(x) \[\large \frac{ 2x^2 + 2x }{ x+1 }\] Factor the top \[\large \frac{ 2x \cancel{(x+1)} }{ \cancel{(x+1)} } = 2x\]

OpenStudy (anonymous):

Ah thanks! I was totally stuck when it comes to actually dividing.